# A Large Scale Analysis of Gender Biases in Text-to-Image Generative Models

### v1

@@ E.1 Retail contexts
In Sec. 4.1, we observe that places in the “retail” cluster are partially strongly female-dominated. Here, we further prove that female-dominated places predominantly relate to fashion, clothes, and beauty. To this end, in Fig. 12, we plot all places in the “retail” cluster where at least one of the five T2I models generated 60% or more female-gendered images. There, we observe that of 14 places, 9 are related to fashion, beauty, or luxury (“beauty salon”, “sweing room”, “dress shop”, “perfume shop”, “wig shop”, “fitting room”, “jewelry shop”, “clothing store”, “fabric store”). In particular, this comprises the most female-dominated retail places.
@@
Further retail places include shopping-related (“drugstore”, “department store”), which we identified as female-associated activity in Sec. 4.1, and “florist shop”, which relates to flowers being a female-leaning type of object.
@@
E.2 Music Instruments
In Sec. 4.2, we find that music instruments make up a male-dominated cluster. This is surprising, as previous research found clear gender associations with respect to musical instruments. In Fig. 13, we show the ratios of female instruments for all objects in the “music instruments” cluster. Note that the objects disc and slipknot are not musical instruments, but we show them nonetheless because they are included in the cluster. The relatively most female-leaning instruments are flute and violin, in accordance with [4, 3]. The same is true for drum, saxophone and guitar, which are male-leaning. However, as also noted in Fig. 5, overall musical instruments are male-leaning and do not follow the associations made by humans.
@@ E.2 Music Instruments
E.3 Dental Jobs
In Sec. 4.2, we take a closer look at the four occupations clustered as “dental jobs”. In 3 of the 4 occupations, the majority of the workforce in the U.S. is women ($>90\%$ for dental hygienist and dental assistant, and $\approx 60\%$ for dental technician) [64]. $\approx 40\%$ of dentists are women. These patterns are reflected in the ratios of female-gendered images generated by T2I models, as shown in Fig. 14. However, SD-3.5-Medium and SD-3-Medium are significantly more biased towards generating female-gendered images than other models.
@@ E.3 Dental Jobs
Appendix F Special Topics
F.1 Activity Classification
For our analyses in Sec. 4.3, we classify our 1405 activities by an LLM to determine if they relate to household chores and work/money. To conduct the classification, we compare 4 different LLMs, namely Llama-3.3-70B-Instruct [27], Qwen2.5-72B-Instruct [97], Yi-1.5-34B [98], and Phi-4 [1]. For classification into household chores, we use the following prompt:
Is the following activity considered a household chore: {activity}. Answer yes or no.
@@ F.1 Activity Classification
In Tab. 6, we show the number of activities that are classified as being related to the two categories, i.e. where the model answers “yes”. Phi-4 labels the fewest activities as household-related or work/money-related, and thus, we proceed with this model, as a lower number of activities makes the analysis more comprehensive. A manual analysis also suggests that the precision of Phi-4 is better than the precision of other models.
@@
F.2 Work-related contexts
@@
where we replace {place} with the context to be classified. We then cluster the resulting 156 work-related places with the method described in Sec. 3.1 and obtain 24 clusters. These are shown in Fig. 15 together with the respective per-model female ratios.
We notice that most clusters are male-dominated, in line with our findings in Sec. 4.1 and Sec. 4.2. The male dominance is particularly strong in clusters related to transportation (shipping) and industrial sites (factories, power/gas/recycling, mine/excavation, …). Female ratios are comparatively higher in contexts related to art (art/television), shopping (market/shop, fashion shops), and places of pleasure (hotel/casino). This confirms our observation that T2I models reflect gender stereotypes and associated work, especially physical labor, with men and social places with women.
@@
We show the resulting 14 places alongside the per-model female ratios in Fig. 16. There, we find that most office-related places are male-dominated. Generally, SD models have higher female ratios across all places. Places with comparatively high female ratios are “call center” and “reception”, which are related to professions where the majority of the workforce are women: Call center employees are listed under “Customer Service Representatives” by [64], and the ratio of women in the U.S. is $65.2\%$. Also, 89.1% of receptionists are women. In SD models, “breakroom” and “office cubicles” are also gender-balanced. In conclusion, the closer analysis of office-related places further strengthens the impression that T2I models associate work more with men than with women.
@@
F.3 Bias Amplification in Occupations
Here, we analyze the relationship between gender ratios in images generated by T2I models and the actual representation of women in the U.S. workforce, as reported by [64]. Of the 575 occupations in our study, [64] provides the percentage of women for 365 occupations. For each occupation prompt $p$, we compute
@@ F.3 Bias Amplification in Occupations
To further explore this perspective, we analyze bias amplification in occupations based on whether the majority of the workforce is male or female. This analysis is presented in Tab. 7. First, we classify each occupation as either “male majority” or “female majority” based on the actual proportion of women in that occupation. If more than 50% of the workforce is female, the occupation is labeled as “female majority”; otherwise, it is labeled as “male majority”. Next, we examine whether the proportion of men or women increases or decreases in images generated by T2I models. If the ratio of the majority group increases, we say the bias is amplified, whereas if it decreases, we say the bias is reduced.
From Tab. 7, we observe that bias in male-majority occupations is almost always amplified. For Flux models, bias in female-majority occupations is more often amplified than reduced. In contrast, for Stable Diffusion models, bias in female-majority occupations is more often reduced than amplified. Overall, these findings confirm our observation that T2I models tend to increase the proportion of men in generated images, while also showing numerous cases where female-majority bias is amplified.
@@
Flux 7.12% 44.11% 19.18% 29.59%
Flux-Schnell 6.03% 45.21% 17.81% 30.96%
SD-3.5-Large 7.67% 43.56% 23.01% 25.75%
SD-3.5-Medium 10.68% 40.55% 28.22% 20.55%
SD-3-Medium 9.32% 41.92% 28.77% 20.00%

### v2

@@ H.1 Retail contexts
In Section 4.1, we observe that places in the “retail” cluster are partially strongly female-dominated. Here, we further prove that female-dominated places predominantly relate to fashion, clothes, and beauty. To this end, in Fig. 11, we plot all places in the “retail” cluster where at least one of the five T2I models generated 60% or more female-gendered images. There, we observe that of 14 places, 9 are related to fashion, beauty, or luxury (“beauty salon”, “sweing room”, “dress shop”, “perfume shop”, “wig shop”, “fitting room”, “jewelry shop”, “clothing store”, “fabric store”). In particular, this comprises the most female-dominated retail places.
@@
Further retail places include shopping-related (“drugstore”, “department store”), which we identified as female-associated activity in Section 4.1, and “florist shop”, which relates to flowers being a female-leaning type of object.
@@
H.2 Music Instruments
In Section 4.2, we find that music instruments make up a male-dominated cluster. This is surprising, as previous research found clear gender associations with respect to musical instruments. In Fig. 12, we show the ratios of female instruments for all objects in the “music instruments” cluster. Note that the objects disc and slipknot are not musical instruments, but we show them nonetheless because they are included in the cluster. The relatively most female-leaning instruments are flute and violin, in accordance with (Abeles & Porter, 1978; Abeles, 2009). The same is true for drum, saxophone and guitar, which are male-leaning. However, as also noted in Fig. 4, overall musical instruments are male-leaning and do not follow the associations made by humans.
@@ H.2 Music Instruments
H.3 Dental Jobs
In Section 4.2, we take a closer look at the four occupations clustered as “dental jobs”. In 3 of the 4 occupations, the majority of the workforce in the U.S. is women ($90\%$ for dental hygienist and dental assistant, and $\approx 60\%$ for dental technician) (Bureau of Labour Statistics, 2023). $\approx 40\%$ of dentists are women. These patterns are reflected in the ratios of female-gendered images generated by T2I models, as shown in Fig. 13. However, SD-3.5-Medium and SD-3-Medium are significantly more biased towards generating female-gendered images than other models.
@@ H.3 Dental Jobs
Appendix I Special Topics
I.1 Work and Money-Making
Figure 14: Work and money-making related clusters of activity prompts.
To assess gender bias regarding work or money-making-related activities, we also classify all 1405 activities by Phi-4 (see Section I.2) and cluster the resulting 139 prompts. This results in 20 clusters, which we label manually and show in Fig. 14.
No cluster other than teaching ($\hat{\mathcal{R}}_{f}\approx 63\%$), work with animals ($\hat{\mathcal{R}}_{f}\approx 62\%$), and writing ($\hat{\mathcal{R}}_{f}\approx 59\%$) contains a majority of female-dominated activities. The cluster with the highest ratio of female-gendered images is teaching, which reflects our previous finding that teachers are associated with women. As already seen in Section 4.1, pet-related activities are frequently associated with women, and linking women with writing resembles the finding that humans associate women more than men with arts (Nosek et al., 2002; Caliskan et al., 2017). Most other money-making activities, including regular work ($\hat{\mathcal{R}}_{f}\approx 34\%$) and money-making ($\hat{\mathcal{R}}_{f}\approx 24\%$), which refers to general activities related to money such as “worrying about money and time”, are male-dominated. We only see higher female ratios for job-seeking activities, i.e. job application ($\hat{\mathcal{R}}_{f}\approx 46\%$) and job interview ($\hat{\mathcal{R}}_{f}\approx 41\%$). This is concerning as underrepresenting women in work- and business-related contexts could reinforce existing stereotypes about women’s role in the workforce, perpetuating or even amplifying limiting gender norms of women as caretakers and men as breadwinners.
I.2 Activity Classification
For our analyses in Section 4.3 and Section I.1, we classify our 1405 activities by an LLM to determine if they relate to household chores and work/money. To conduct the classification, we compare 4 different LLMs, namely Llama-3.3-70B-Instruct (Dubey et al., 2024), Qwen2.5-72B-Instruct (Yang et al., 2024), Yi-1.5-34B (Young et al., 2024), and Phi-4 (Abdin et al., 2024). For classification into household chores, we use the following prompt:
Is the following activity considered a household chore: {activity}. Answer yes or no
@@ I.2 Activity Classification
In Table 7, we show the number of activities that are classified as being related to the two categories, i.e. where the model answers “yes”. Phi-4 labels the fewest activities as household-related or work/money-related, and thus, we proceed with this model, as a lower number of activities makes the analysis more comprehensive. A manual analysis also suggests that the precision of Phi-4 is better than the precision of other models.
@@
I.3 Work-related contexts
@@
where we replace {place} with the context to be classified. We then cluster the resulting 156 work-related places with the method described in Section 3.1 and obtain 24 clusters. These are shown in Fig. 15 together with the respective per-model female ratios.
We notice that most clusters are male-dominated, in line with our findings in Section 4.1 and Section 4.2. The male dominance is particularly strong in clusters related to transportation (shipping) and industrial sites (factories, power/gas/recycling, mine/excavation, …). Female ratios are comparatively higher in contexts related to art (art/television), shopping (market/shop, fashion shops), and places of pleasure (hotel/casino). This confirms our observation that T2I models reflect gender stereotypes and associated work, especially physical labor, with men and social places with women.
@@
We show the resulting 14 places alongside the per-model female ratios in Fig. 16. There, we find that most office-related places are male-dominated. Generally, SD models have higher female ratios across all places. Places with comparatively high female ratios are “call center” and “reception”, which are related to professions where the majority of the workforce are women: Call center employees are listed under “Customer Service Representatives” by Bureau of Labour Statistics (2023), and the ratio of women in the U.S. is $65.2\%$. Also, 89.1% of receptionists are women. In SD models, “breakroom” and “office cubicles” are also gender-balanced. In conclusion, the closer analysis of office-related places further strengthens the impression that T2I models associate work more with men than with women.
@@
I.4 Bias Amplification in Activities
To analyze bias amplification in activities, we retrieve images from the LAION-400m dataset (Schuhmann et al., 2021) that match our activity prompts. We chose LAION-400m because it is representative of the web-scale datasets typically used to train T2I image models. To avoid biases in CLIP-based retrieval (Hirota et al., 2025; Berg et al., 2022; Seth et al., 2023), we use a text-based retrieval method: using spaCy, we extract all non-stopword lemmas from both activity prompts and captions in LAION-400m. We match a prompt to a caption if all the prompt’s lemmas are contained in the caption’s lemmas, i.e. if the prompt lemmas are a subset of the caption lemmas. If more than 10,000 images match a single activity prompt, we randomly sample 10,000 images for further analysis.
For each matched image, we detect people bounding boxes by YOLOv10 and assign perceived gender using InternVL2-8B, with the same prompt setup described in Section 3.2 and Section E.1. We apply the same filtering to LAION-400m images as we do to images generated by T2I models: we discard any image with no recognizable gender or with both men and women present. After filtering, 152 activity prompts remain, each with at least 50 matched LAION-400m images. We use these to estimate the proportion of female-gendered images for each activity in LAION-400m. The average female ratio across these 152 activity prompts is approximately $52\%$, suggesting that this subset of activities is not strongly biased toward either gender. In contrast, the average female ratio in generated images is only around $41\%$, indicating that women are underrepresented in generated images even when compared to web-scale data.
To analyze this more closely, we categorize activities as either “female majority” (activities where more than 50% of the LAION-400m images are female-gendered) or “male majority” (where more than 50% are male-gendered). For each activity, we then check whether the ratio of the majority gender increases or decreases in images generated by T2I models. If the ratio increases, we call it bias amplification, and if it decreases, we call it bias reduction.
Detailed results are shown in Table 8. We find that male-majority activities tend to show an even higher male ratio in generated images. For female-majority activities, the outcomes are more balanced between amplification and reduction. However, overall, female-majority activities tend to have a lower female ratio in generated images than in LAION-400m. This suggests that models amplify gender imbalances in favor of male-gendered images, even beyond what is present in the pretraining data, and this applies to categories beyond occupations. To fully understand the causes of these effects, a more detailed analysis of web-scale image datasets is needed; for example, the overall ratio of men and women in the pretraining data remains unknown.
Male majority Female majority
reduced amplified reduced amplified
Flux 12.68% 87.32% 60.49% 39.51%
Flux-Schnell 18.31% 81.69% 71.60% 28.40%
SD-3.5-Large 25.35% 74.65% 60.49% 39.51%
SD-3.5-Medium 35.21% 64.79% 40.74% 59.26%
SD-3-Medium 16.90% 83.10% 60.49% 39.51%
Table 8: We classify activities into “male majority” or “female majority” based on whether there are more male-gendered images than female-gendered images in LAION-400m. Then, we check if, in generated images, the majority gender has increased or decreased ratio. If the majority gender is increased, we label it as “amplified”; if the majority gender is decreased, we label the occupation as “reduced”.
I.5 Bias Amplification in Occupations
Here, we analyze the relationship between gender ratios in images generated by T2I models and the actual representation of women in the U.S. workforce, as reported by Bureau of Labour Statistics (2023). Of the 575 occupations in our study, Bureau of Labour Statistics (2023) provides the percentage of women for 365 occupations. For each occupation prompt $p$, we compute
@@ I.5 Bias Amplification in Occupations
To further explore this perspective, we analyze bias amplification in occupations based on whether the majority of the workforce is male or female. This analysis is presented in Table 9. First, we classify each occupation as either “male majority” or “female majority” based on the actual proportion of women in that occupation. If more than 50% of the workforce is female, the occupation is labeled as “female majority”; otherwise, it is labeled as “male majority”. Next, we examine whether the proportion of men or women increases or decreases in images generated by T2I models. If the ratio of the majority group increases, we say the bias is amplified, whereas if it decreases, we say the bias is reduced.
From Table 9, we observe that bias in male-majority occupations is almost always amplified. For Flux models, bias in female-majority occupations is more often amplified than reduced. In contrast, for Stable Diffusion models, bias in female-majority occupations is more often reduced than amplified. Overall, these findings confirm our observation that T2I models tend to increase the proportion of men in generated images, while also showing numerous cases where female-majority bias is amplified.
@@
Flux 6.85% 44.38% 19.18% 29.59%
Flux-Schnell 6.30% 44.93% 20.00% 28.77%
SD-3.5-Large 7.67% 43.56% 23.56% 25.21%
SD-3.5-Medium 10.96% 40.27% 28.77% 20.00%
SD-3-Medium 8.77% 42.47% 29.32% 19.45%
@@
Appendix J Detailed Discussion of Limitations
White Black East Asian Latino-Hispanic Middle Eastern Indian SE Asian Other
a trout 0.99 going to a reggae concert 0.87 studying mandarin chinese. 0.99 eating tacos. 0.36 praying the obligatory 5 daily prayers. 0.59 in the slum 0.33 near the rice paddy 0.67 playing skyrim 0.34
going to a bass pro elite competition. 0.99 a basketball 0.67 watching anime 0.95 buying an awesome burrito. 0.32 in the medina 0.49 inside the kasbah 0.31 in the slum 0.34 playing world of warcraft. 0.33
by the fjord 0.99 on the savanna 0.61 a china 0.95 fast food worker 0.28 at the caravansary 0.49 at the temple 0.25 agricultural worker 0.30 playing the computer game lords of the fallen. 0.11
reading the new anthology with christine feehan in it. 0.99 tutoring their basketball players before their history exam. 0.59 brewing tea gong-fu style. 0.94 making guacamole 0.27 outside the mosque 0.47 in the village 0.24 in the village 0.28 a spear 0.07
flying to the adirondacks with their girlfriend or boyfriend. 0.99 usher 0.59 in the japanese garden 0.91 a bikini 0.25 inside the kasbah 0.44 inside the fort 0.24 at the temple 0.27 the hoodoo 0.06
trimming their beard. 0.98 at the basketball court 0.58 going out to dinner with their family to enjoy delicious chinese food. 0.90 eating a burrito bowl at chipotle. 0.24 going to a far away place for religious reasons. 0.38 laborer 0.20 farming or fishing worker 0.27 watching game of thrones 0.04
at the hunting lodge 0.98 playing basketball 0.56 a japan 0.90 licensed practical nurse 0.24 religious worker 0.37 in the medina 0.19 near the garbage dump 0.25 playing mario 0.03
installing a new door sweep. 0.98 meeting up with a friend and playing basketball for the entire afternoon. 0.54 going to hunan garden with their girlfriend/boyfriend. 0.85 registered nurse 0.24 chef 0.32 an elephant 0.18 rice 0.22 watching all of the lord of the rings movies. 0.03
hiking with their dog. 0.98 preparing for the upcoming fantasy football draft. 0.51 reading a book called ”the taker” by alma katsu. 0.84 meat processing worker 0.23 near the mastaba 0.30 near the garbage dump 0.16 at the bazaar 0.19 a squid 0.03
hiking 0.98 working on their fantasy football lineup. 0.47 in the zen garden 0.84 physician assistant 0.23 baker 0.25 at the bazaar 0.15 within the rainforest 0.18 within the rainforest 0.03
Table 10: Top 10 prompts with highest avg. ratio of generated people for each race across T2I models.
While our study makes a valuable contribution to understanding gender bias in current T2I models and extends insights from previous work, there are several important areas that we do not address. These include gender identities beyond the binary, social categories beyond gender, intersectional biases, and debiasing techniques. Below, we explain why these topics cannot currently be properly analyzed using the methods applied in our study. Furthermore, we justify our use of automatic methods for labeling perceived gender.
Non-binary gender identities. In the generated images, we do not find clear evidence of images that unambiguously depict non-binary gender identities. We believe that such an analysis should involve judgments or annotations from people who identify as non-binary, similar to (Ungless et al., 2023). Without this input, it is unclear how to identify relevant images or analyze stereotypes within them. This is also supported by our findings in Section E.2. Currently, automatic methods do not label images with “nonbinary”, and as mentioned above, we are not aware of any other techniques that enable automatic analysis of images that may depict non-binary gender identities.
Automatic gender labeling. Using automatic methods to assign sensitive attributes such as gender (as well as race or age) can be problematic because models may introduce errors, carry their own biases, and in doing so, undermine the validity of analyses based on automatic labels. Even worse, if models are biased, they may reinforce those biases throughout the analysis. At the same time, using automatic tools is essential for conducting large-scale studies like ours. Therefore, we take steps to ensure our results are as valid as possible by addressing issues that arise from automatic methods. First, we filter images based on detected people, using state-of-the-art object detectors (Wang et al., 2024). Then, we crop person bounding boxes to reduce bias from background or contextual elements. Most importantly, we evaluate whether gender assignments from InternVL2-8B align with human annotations of perceived gender. As shown in Section E.1, this is indeed the case. Given the near-perfect alignment between human labels and automatically determined labels, we do not expect automatic methods to introduce significantly more errors or reinforce stereotypes beyond what human annotators might. While gender bias remains a concern in MLLMs (Girrbach et al., 2025), it is less pronounced in discriminative tasks that aim specifically to label gender.
Debiasing methods. The aim of our study is to provide a detailed, in-depth analysis of gender bias in current T2I models across everyday scenarios. In addition to understanding the societal issues related to T2I models, exploring ways to address these problems is also an important area of research. However, as models continue to be used without explicit steering mechanisms (Zhang et al., 2023; De Simone et al., 2023; Clemmer et al., 2024; Brack et al., 2023), it becomes crucial to develop a clear understanding of their underlying issues. Determining how and when to apply steering or other debiasing techniques is another complex challenge, which lies beyond the scope of this study. For instance, it remains an open question whether solutions to these identified problems should be implemented by model providers or users. One possible approach is “ambiguity in, diversity out” (Katirai et al., 2024), although this too raises concerns, such as maintaining contextual appropriateness. Given these challenges, detailed insights into model biases, like those provided in our study, are essential for making informed decisions about modifying or restricting model outputs. For the same reason, we do not aim to develop a benchmark. The fact that models exhibit bias has been shown before, and benchmarks typically construct one or a few measures of bias that help guide researchers and developers toward creating less biased models. However, such benchmarks can only indicate the degree of bias, not the specific manifestations of bias that we provide in this study.
Nonetheless, our findings are important and will inform future development of debiasing strategies. The main areas where our findings are relevant are Targeted Data Curation, as our fine-grained analysis identifies specific concepts and contexts (e.g., “laundry”, occupations involving physical work) where gender bias is most severe. This can inform mitigation efforts like targeted data collection or strategic rebalancing of existing training sets. Also, our work informs Hybrid Mitigation Strategies targeting data and models alike. Here, our finding that models can amplify bias present in web-scale data (LAION-400M) is a crucial insight, because it implies that data-centric approaches alone may be insufficient. Future work should investigate hybrid strategies that address bias in both the training data and the training/inference process of the model.
Social categories beyond gender. We find that T2I models show strong biases in other social categories, such as race and age, when generating images from the underspecified prompts used in our study. To illustrate this, we detect perceived race and age for all identified people in the generated images using InternVL2-8B. The prompts used are similar to those employed for detecting perceived gender. For detecting race and age, we use the following prompts:
What is the race of the person in the image?
A. black
B. east asian
C. indian
D. middle eastern
E. latino-hispanic
F. southeast asian
G. white
H. other
I. unclear/cannot tell
Answer with a letter (A, B, C, etc.).
And for age, we use the following prompt:
What is the age of the person in the image?
A. less than 20
B. 20-29
C. 30-39
D. 40-49
E. 50-59
F. 60-69
G. more than 70
H. unclear/cannot tell
Answer with a letter (A, B, C, etc.).
In both cases, we randomly permute the option order (but not the option letters) to avoid label bias. Race and age categories are from FairFace (Karkkainen & Joo, 2021). However, we truncate the underage age categories to a single label (“less than 20”).
We then calculate the overall ratios of people assigned to each race and age category for all 5 models in this study. Before calculating ratios, we drop all people who receive the “unclear/cannot tell” label. Results for race are in Fig. 18 and for age in Fig. 19. From these results, it is clear that models predominantly generate white and young (age 20-29 or 30-39) people, confirming results in (Wu et al., 2024; Ghosh & Caliskan, 2023).
Figure 18: Race ratios for people in all images generated by T2I models, as assigned by InternVL2-8B.
Figure 19: Age ratios for people in all images generated by T2I models, as assigned by InternVL2-8B.
In Table 10, we also show the top 10 prompts with the highest average ratio of generated people across models for each race. There, we find that White and East Asian individuals have a notable number of prompts that, consistently across models, generate predominantly images of the respective race in all T2I models. Moreover, only prompts associated with White people tend to be fairly general, while prompts linked to other races are mostly tied to cultural or national stereotypes. For example, East Asian-looking people are generated from prompts mentioning East Asian cultural elements, such as “anime” or “mandarin chinese”, while Latino-looking people appear in images generated from prompts like “tacos” or “burrito”. An analysis of such cultural stereotypes in T2I models has been conducted in (Dehdashtian et al., 2025; Jha et al., 2024).
Based on these findings, we conclude that the dominance of young, White individuals in generated images makes it difficult to perform intersectional analysis under the current experimental settings. To properly study race and age biases, as well as their intersection, it is necessary to explicitly prompt T2I models for these attributes and analyze the resulting stereotypes.
Lastly, we validate the performance of MLLM race and age detection using human annotations from the FairFace dataset. Using the prompts described above, we assign race and age labels to all images in the FairFace validation set. Detailed results are shown in Table 11. Overall, the accuracy is 68% for race and 56% for age. While these values are lower than the reported accuracies for gender, they are still significantly better than random chance, considering the larger number of categories. Therefore, we conclude that our observations about race and age stereotypes are approximately accurate, although a fine-grained analysis remains difficult due to the lower agreement between automatic methods and human labels.
Precision Recall F1-Score Support
Black 0.83 0.91 0.87 1556
East Asian 0.67 0.86 0.75 1550
Indian 0.83 0.65 0.73 1516
Latino-Hispanic 0.56 0.53 0.55 1623
Middle Eastern 0.62 0.54 0.57 1209
Southeast Asian 0.58 0.48 0.53 1415
White 0.75 0.74 0.74 2085
(a) Detailed race labeling results by InternVL2-8B wrt. human annotations on the FairFace validation set. Overall accuracy is 68%.
Precision Recall F1-Score Support
20-29 0.66 0.60 0.63 3300
30-39 0.48 0.45 0.47 2330
40-49 0.48 0.22 0.30 1353
50-59 0.42 0.37 0.39 796
60-69 0.30 0.56 0.39 321
less than 20 0.81 0.82 0.81 2736
more than 70 0.34 0.60 0.43 118
(b) Detailed age labeling results by InternVL2-8B wrt. human annotations on the FairFace validation set. Overall ccuracy is 56%.
Table 11: Race and age classification results on the FairFace validation set.
Appendix K Detailed Comparison to Previous Work
In this section, we provide a detailed comparison between our work and previous studies on analyzing gender bias in T2I models. We focus on works that use gender-neutral prompts, as this matches the experimental setup in our study. The comparison is shown in LABEL:tab:supp:detailedcomparisontoprevious. For each paper, we include the number of gender-neutral prompts, the total number of images generated per evaluated model, a brief summary of the main findings, and a short note on how our study differs from that work.
In comparison to previous work, our study significantly improves the understanding of gender bias in T2I models by offering a detailed analysis across a wide range of everyday activities, places, objects, and occupations. As noted by Wan et al. (2024) and clearly shown in LABEL:tab:supp:detailedcomparisontoprevious, most prior studies have focused mainly on occupational prompts to highlight bias. While this focus is valuable, examining gender bias beyond occupations is also essential for a more complete understanding of how such bias manifests in T2I models.
Another aspect is the typically very small scale of studies, as also shown in LABEL:tab:supp:detailedcomparisontoprevious. While this allows us to conclude that models are biased, gaining concrete insights into these biases requires a broader analysis like ours. Two other studies also generate a large number of images: Wu et al. (2024) generated images from 200,000 distinct prompts, but used them not to analyze gender distributions for individual prompts or prompt groups, but to examine representational similarities between images from gender-neutral and gendered prompts. This setup is well-suited to reveal an overall male bias in the evaluated models, but does not support a detailed analysis of the specific stereotypes replicated by the models. Similarly, Luo et al. (2024; 2025) (these two papers have significant textual overlap) generated images for 2,123,200 prompts, about 70% of which focus on occupations. This study uses the images to compute holistic bias scores for comparing and ranking models, whereas our goal is to document biases in detail.
Table 12: Detailed comparison to previous work. We show the number of images in the study for each evaluated model, the number of prompts, a summary of the study’s findings, and a comment on how our study contributes beyond the respective prior work.
# Images # Prompts Main Findings Novelty of our work
Bianchi et al. (2023) 2000 20 This study includes 20 gender-neutral prompts (in addition to prompts that either explicitly specify gender and race or focus on objects from diverse cultural contexts). Of these, 10 prompts describe people (e.g., “an exotic person”, “a terrorist”), and 10 describe occupations. The paper reports on how gender and racial stereotypes are reflected and perpetuated in these 20 analyzed cases. Our study enables a more thorough analysis of gender bias in T2I image models by including a larger set of prompts and generated images. This makes it possible to automatically evaluate broader trends, such as associations with household chores or workplaces, beyond just a few manually examined examples.
Cheong et al. (2024) 1050 105 Generated 10 images each for 105 different occupations. After collecting gender and race annotations from human labelers, a filtered set of 67 occupations was compared with respect to race and gender ratios in the U.S. workforce and the generated images. The study finds strong bias amplification, i.e., the images often depict only men or only women for a given occupation. Our study examines gender bias not only in occupations but also in related categories such as activities, places, and objects. Within occupations, we include the complete set from the U.S. BLS list. Our method produces more reliable estimates of gender ratios by sampling a larger number of images and filtering out unsuitable prompts and images.
Chinchure et al. (2024) 5328 111 This study proposes a method to detect biases in generated images on dynamically identified bias axes. First, an LLM proposes potentially relevant bias axes from the given prompt. Then, after a set of images has been generated from the prompt, these biases are verified or rejected based on counterfactual image sets and VQA attribute detections. The relevance of the detected biases is verified through human judgments. In contrast to proposing a general method to discover various biases, we specifically analyze gender bias regarding a broad range of activities, contexts, objects, and occupations. We see our research and the method proposed in Chinchure et al. (2024) as orthogonal and mutually beneficial: Once systematic biases in T2I models have been detected through methods like TIBET, large-scale analyses such as those conducted in our study will give a deep understanding of how exactly these detected biases surface.
Cho et al. (2023) 747 83 Generated 9 images based on 83 gender-neutral occupation prompts (excluding variants that explicitly specify gender). Gender, skin tone, and 15 other attributes were automatically detected. The results show that T2I models generally generate more men than women. Additionally, skirts appear only on women, while suits are more commonly shown on men. Our study analyzes gender bias not only in occupation-related prompts but also in everyday activities and locations. In addition, our evaluation protocol provides a more reliable estimate of gender ratios by sampling more images and filtering out unsuitable ones. Lastly, we reduce contextual bias in automatic gender detection by cropping the images to focus on person bounding boxes.
Luccioni et al. (2024) 4380 146 Generated 30 images using 146 gender-neutral occupation prompts. Gender and race distributions were analyzed with a non-parametric method that does not rely on explicit gender or race labels. A comparison with U.S. BLS statistics shows that women — especially Black women — are underrepresented. We analyze gender biases beyond just occupations while also including a larger set of occupations. This broader analysis helps us identify bias trends on a wider scale. At the same time, we ensure our results are reliable by using large-scale sampling and filtering.
Luo et al. (2024; 2025) 2 123 200 2654 Generated 2,654 prompts related to occupations, social relationships, and attributes. A significant portion of the prompts include explicit gender or race identifiers, and about 70% of the prompts focus on occupations. The study evaluates different models using various bias scores to determine which ones are the most or least biased. While this study also provides a large-scale evaluation of bias in T2I models, our work offers two key contributions: First, instead of presenting overall bias scores to compare models, we closely examine which specific biases (e.g., those related to household chores) the models exhibit. Second, our study goes beyond occupations, which are the main focus of the FaintBench benchmark, and explores a broader range of categories.
Lyu et al. (2025) 2000 100 This study evaluates the reliability and validity of gender bias analysis pipelines, identifying several issues, such as images featuring people of different genders or no people at all. The prompts used in the analysis cover all the categories included in this study, but on a much smaller scale (10 to 40 prompts). Our study focuses on a detailed analysis of gender bias in T2I models at a large scale. To achieve this, we include a significantly higher number of prompts and analyses, comparing our results to those related to human stereotypes. However, the insights from this study shaped our experimental design, particularly emphasizing the need for careful and rigorous filtering.
Seshadri et al. (2024) 31000 62 This study investigates bias amplification using 62 occupation prompts and concludes that bias amplification is largely explained by distribution shifts between the training and probing distributions. Our study thoroughly documents the gender bias in recent T2I models, including observations of bias amplification. However, we do not explore the causes behind this bias amplification. Instead, we analyze a broad range of activities, places, objects, and occupations to provide in-depth insights.
Ungless et al. (2023) 924 231 Generated 4 images for each of 321 prompts centered on non-binary identities. The key findings are that non-binary identities are poorly represented by T2I models, often resulting in the creation of NSFW or degrading content. Our study focuses specifically on binary gender. We also note that, without explicit instructions, models do not produce images that clearly represent non-binary identities. As a result, it is currently impossible to quantitatively explore biases related to non-binary identities using the models and methods applied in this study. However, we believe that addressing this issue is an important direction for future research.
Wu et al. (2024) 800 000 200 000 This study examines how gender-neutral prompts are represented across different T2I models (text, latent noise, and images). The key finding is that when gender (or other image characteristics) is not specified in the prompt, the generated images tend to resemble those created from masculine prompts. While this study analyzes a large number of prompts, it does not estimate gender ratios for prompts or prompt groups. Instead, it focuses on examining representational similarities. In contrast, our method allows for a deeper exploration of biases across a wide range of activities, places, objects, and occupations. This approach enables us to make precise statements about whether models display specific types of gender bias.